OpenStudy (anonymous):

Find an equation, in slope-intercept form, that passes through point (-2,1) with slope 2

OpenStudy (anonymous):

Can Someone Please Tell me the answer for this problem?

sam (.sam.):

\[y_{1}=1~~,~~x_{1}=-2~~,~~m=2\] Use the formula \[y-y_{1}=m(x-x_{1})\] Substitute and get the equation

OpenStudy (anonymous):

i already have the answer im just trying to get someone to tell me an answer to see if i got it right

OpenStudy (anonymous):

y = 2x + 5

sam (.sam.):

y-1=2(x--2) y=2x+4+1 y=2x+5

sam (.sam.):

yep

OpenStudy (anonymous):

that my answer i dont know if im correct

sam (.sam.):

correct

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

Hey Is This TRUE or FALSE??The slope-intercept form of the equation of the line with slope 0 passing through the point (-2,7) is y = 7.